How many fourths are there in 3?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Here, we will proceed by discussing about one by fourth of any number and finding the value of one by fourth of number 3. Then, we will use the formula i.e., Number of fourths in any number = Number(14). Complete step-by-step solution - As we know that one by fourth of any number can be obtained by dividing that number with four. The one by fourth of any number represents that when we will be combining four one by fourth of that number, we will be getting that number. In the physical world, the interpretation of one by fourth of any item refers to dividing that item into four parts and each part of this item will represent one by fourth of the original item. For number = 3, One by fourth of 3 = 34 In order to find out the total number of fourths which will be there in any number, we will divide this number with the one by fourth of that number. i.e., Number of fourths in any number = Number(14) →(1) Using equation (1), we get Number of fourths of 3 = 3(14)=3×41=12 Therefore, there are a total 12 fourths present in 3. Note- The expression for one by fourth of any number is given by One by fourth of any number = Number4. Here, if instead of asking for the number of fourths which are there in 3 we are asked for the number of threes which are there in 3. For this we will be using the formula i.e., Number of threes in any number = Number(13).